A toy rocket is propelled upward at a speed of 96ft/sec by a machine that is 11 feet off the ground. The height h of the object after t seconds can be found using the function: h (t) = -16t^2 +96t + 11
What is the height6 of the rocket after 3 seconds and when will the toy rocket reach the height of 122 feet?

Answer by ikleyn(52778)

.

To answer first question, substitute t= 3 seconds into the formula and calculate

    h(3) = -16*3^2 + 96*3 + 11.


It is simple arithmetic.



To answer second question, solve this quadratic equation

    -16t^2 + 96t + 11 = 0.


It is equivalent to equation

    16t^2 - 96t - 11 = 0.


Use the quadratic formula. It is simple arithmetic, too.

Happy calculations (!)

----------------------

In this site,  there is a bunch of lessons on a projectile thrown/shot/launched vertically up

    - Introductory lesson on a projectile thrown-shot-launched vertically up
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.
Read them attentively and learn on how to solve this type of problems once and for all.